Write the slope-intercept equation for the line that passes through (-4, -3) and (-6, 7). Please show all of your work.

Answer by tarungehlot(9)

slope = m= y2- y1/x2-x1= 7+3/-6+4=10/-2=-5
now slope intercept form or two point from
y -y1 = m( x-x1)
y+3= -5 ( x+4)
y+3= -5x -20
y= -5x -23
compare with
y= mx +c
slope m=-5
intercept c=-23